# Scalable Decentralized Algorithms for Online Personalized Mean Estimation

### v3

@@ Abstract
In numerous settings, agents lack sufficient data to directly learn a model. Collaborating with other agents may help, but introduces a bias-variance trade-off when local data distributions differ. A key challenge is for each agent to identify clients with similar distributions while learning the model, a problem that remains largely unresolved. This study focuses on a simplified version of the overarching problem, where each agent collects samples from a real-valued distribution over time to estimate its mean. Existing algorithms face impractical per-agent space and time complexities (linear in the number of agents $|\mathcal{A}|$). To address scalability challenges, we propose a framework where agents self-organize into a graph, allowing each agent to communicate with only a selected number of peers $r$. We introduce two collaborative mean estimation algorithms: one uses a message-passing scheme, while the other employs a consensus-based approach, with complexity $\mathcal{O}(r\cdot\log|\mathcal{A}|)$ and $\mathcal{O}(r)$, respectively. We establish conditions under which both algorithms yield asymptotically optimal estimates and offer a theoretical characterization of their performance.
@@ 1 Introduction
Users’ devices have become increasingly sophisticated and generate vast amounts of data. This wealth of data has paved the way for the development of accurate and complex models. However, it has also introduced challenges related to security, privacy, real-time processing, and resource constraints. In response, Federated Learning (FL) has emerged as a pivotal privacy-preserving methodology for collaborative model training [Kairouz et al., 2021; Li et al., 2020]. Traditional FL methods focus on learning a unified model applicable to all clients. Yet, the statistical diversity observed across clients’ datasets has led to the pursuit of personalized models. These models aim to better align with the unique data distributions of individual clients [e.g., Ghosh et al., 2020; Fallah et al., 2020; Li et al., 2021; Marfoq et al., 2021; Ding and Wang, 2022].
@@
Many personalized FL strategies involve grouping clients into homogeneous clusters to then tailor a model for each cluster [e.g., Ghosh et al., 2020; Sattler et al., 2021; Ding and Wang, 2022]. The optimal clustering would ideally group clients whose local optimal models are similar. However, the inherent challenge lies in the fact that these optimal models are unknown a priori. This dilemma results in the two tasks—model learning for the specified task and cluster identification—being deeply interconnected. Various studies have suggested empirical measures of similarity as a workaround [e.g., Ghosh et al., 2020; Sattler et al., 2021], while others rely on presumed knowledge of distances across data distributions [e.g. Ding and Wang, 2022; Even et al., 2022]. Nonetheless, accurately estimating these distances, especially within a FL framework where clients may possess limited data, proves to be particularly challenging. Such estimation difficulties are well discussed in the literature (e.g., in Even et al. [2022] [Sec. 6]), underscoring the problem of identifying similar clients for collaborative model learning as a significant, yet unresolved issue.
@@
In this paper, we narrow our focus to a fundamental aspect of the broader challenge, specifically aiming to estimate the mean of a real-valued distribution, considered as the archetypal federated learning problem in many papers [Dorner et al., 2024; Tsoy et al., 2024; Grimberg et al., 2021]. This task holds significant practical relevance across various fields, such as smart agriculture, grid management, and healthcare, where multiple sensors collect private, noisy data on identical or related variables Adi et al. [2020]. Importantly, the task of personalized mean estimation serves as a foundational problem that can illuminate key aspects of personalized federated learning.
@@
We are interested in an online decentralized version of personalized mean estimation, where at each time slot, a client collects new samples and can exchange information with a limited number of other clients. To the best of our knowledge, the state-of-the-art method in this setting is the Collaborative Mean Estimation algorithm, ColME, by Asadi et al. [2022]. Unfortunately, ColME struggles with scalability in large systems, as its per-agent space and time complexities are both linear in the number of clients $|\mathcal{A}|$.
@@
To address these scalability challenges, we propose that clients self-organize into a network where each client communicates with at most a fixed set of $r$ neighbors. Over time, this set of neighbors is pruned, as clients progressively exclude those that are more dissimilar. Within this context, we introduce two collaborative mean estimation algorithms: one uses a message-passing scheme and the other is based on consensus. The complexities of these algorithms are $\mathcal{O}(r\cdot\log|\mathcal{A}|)$ and $\mathcal{O}(r)$, respectively. We demonstrate that, even though each client exchanges information with only $r\ll|\mathcal{A}|$ clients, it is possible to achieve with high probability a speedup in the convergence time of mean estimates by a factor of $\Omega(|\mathcal{A}|^{1/2-\phi})$, with $\phi$ arbitrarily close to 00.
@@ 2 Related Work
For an overview of the field of personalized federated learning, we refer the reader to the recent survey Tan et al. [2023]. Here, we limit to mention the most relevant approaches for this paper.
@@
Ghosh et al. [2020] and Sattler et al. [2021] were the first to propose clustered FL algorithms, which split the clients on the basis of the similarity of their data distributions; the similarity is evaluated empirically by the Euclidean distance of the local models and by the cosine similarity of their updates. Ding and Wang [2022] study more sophisticated clustering algorithms assuming that clients can efficiently estimate some specific (pseudo)-distances across local distributions (i.e., the integral probability metrics).
@@
Beaussart et al. [2021], Chayti et al. [2022], Grimberg et al. [2021], and Even et al. [2022] consider decentralized approaches, which allow each client to learn a personal model relying on a specific convex combination of information (gradients) from other clients. In particular, Even et al. [2022] prove that collaboration can at most speed up the convergence time linearly in the number of similar agents and provide algorithms, which, under apriori knowledge of pairwise client distributions’ distances, achieve such speedup. The authors recognize the complexity of estimating these distances and provide practical estimation algorithms for linear regression problems which asymptotically achieve the same speedup scaling the number of clients but maintaining the number of clusters fixed.
The generalization properties of personalized models obtained by convex combinations of clients’ models are studied in Mansour et al. [2020], while Donahue and Kleinberg [2021] look at the problem through the lens of game theory. The work most similar to ours is Asadi et al. [2022], which we describe in detail in the next section.
3 Model and Background on ColME
@@
We consider a set $\mathcal{A}$ of agents (computational units). Each agent $a\in\mathcal{A}=\{1,2,\cdots,|\mathcal{A}|\}$ generates at each time instant $t$ a new sample $x_{a}^{t}\in\mathbb{R}$ drawn i.i.d. from distribution $D_{a}$ with expected value $\mu_{a}=\mathbb{E}[x_{a}^{t}]$. Expected values are not necessarily distinct across agents. Given two agents $a$ and $a^{\prime}$, we denote by $\Delta_{a,a^{\prime}}:=|\mu_{a}-\mu_{a^{\prime}}|$ the gap between the agents’ true means, and with $\mathcal{C}_{a}$, the group of agents with the same true mean of $a$, (i.e. such that $\Delta_{a,a^{\prime}}=0$). In the following we will refer to $\mathcal{C}_{a}$ as the ‘similarity class of $a$.’
@@
The goal of each agent $a\in\mathcal{A}$ is to estimate $\mu_{a}$. To this purpose, the client can compute its local average estimate $\bar{x}^{t}_{a,a}=\frac{1}{t}\sum_{1}^{t}x^{t^{\prime}}_{a}$, but can also obtain a more accurate estimates taking advantage of information from other agents in its similarity class.
Asadi et al. [2022] proposed ColME as a collaborative algorithm for mean estimation. ColME relies on two key steps: i) trying to identify the similarity classes; ii) improving local mean estimates by sharing information with agents believed to belong to the same class.
@@
We denote by $\mathcal{C}_{a}^{t}$ the set of agents that agent $a$ estimates, at time $t$, to be part of its own similarity class $\mathcal{C}_{a}$. Agent $a$ decides that agent $a^{\prime}$ belongs to the same class if their local average estimates are sufficiently close, specifically, if two appropriately defined confidence intervals centered on these estimates overlap. In general, at time $t$, agent $a$ will not have access to the most recent local average estimates of agent $a^{\prime}$, but to a stale value, corresponding to the last time the two clients have exchanged information. We denote such value as $\bar{x}_{a,a^{\prime}}^{t}$ and the corresponding number of samples as $n_{a,a^{\prime}}^{t}$ [^1] Note that $n_{a,a^{\prime}}^{t}\leqslant t$ coincides with the last time-slot at which the two agents have communicated.. Agent $a$ can then estimate its true mean and the true mean of agent $a^{\prime}$ to belong to the confidence intervals $I_{a,a}=\left[\bar{x}^{t}_{a,a}-\beta_{\gamma}(t),\bar{x}^{t}_{a,a}+\beta_{\gamma}(t)\right]$, $I_{a,a^{\prime}}~{}=~{}\left[\bar{x}^{t}_{a,a^{\prime}}-\beta_{\gamma}(n^{t}_{a,a^{\prime}}),\bar{x}^{t}_{a,a^{\prime}}+\beta_{\gamma}(n^{t}_{a,a^{\prime}})\right]$, respectively. As expected, the interval amplitude $\beta_{\gamma}(n)$ depends on the number of samples $n$ on which the empirical average is computed, and on the target level of confidence $1-2\gamma$ associated with the interval. Agent $a$ will then assume that $a^{\prime}$ belongs to its similarity class $\mathcal{C}_{a}$ if the two intervals overlap, i.e., $I_{a,a}\cap I_{a,a^{\prime}}\neq\emptyset$, or equivalently if:
$d^{t}_{\gamma}(a,a^{\prime}):=|\bar{x}_{a,a}-\bar{x}_{a,a^{\prime}}|-\beta_{\gamma}(t)-\beta_{\gamma}(n^{t}_{a,a^{\prime}})<0.$ (1)
In order to estimate its similarity class in the most accurate way, agent $a$ could retrieve at eath time $t$ the most recent estimate $\bar{x}^{t}_{a^{\prime},a^{\prime}}$ from each other agent $a^{\prime}$. This, however, would lead to a quadratic communication burden $\Theta(\mathcal{|A|}^{2})$. To limit such a burden, node $a$ cyclically queries a single node in $\mathcal{C}_{a}^{t-1}$ at each time instant $t$, according to a Round-Robin scheme. This leads to the following update rule for $\mathcal{C}_{a}^{t}$:
@@
and we observe that by construction $\mathcal{C}_{a}^{t}\subseteq\mathcal{C}_{a}^{t-1}$. Initially, the agent sets $\mathcal{C}_{a}^{0}=\mathcal{A}$ and progressively takes irreversible decisions to remove agents.
@@
Each node $a$ computes an estimate $\hat{\mu}_{a}^{t}$ of its true mean $\mu_{a}$ by combining the available estimates according to a simple weighting scheme, where number of samples are the weights:
$\hat{\mu}_{a}^{t}=\sum_{a^{\prime}\in\mathcal{C}_{a}^{t}}\frac{n^{t}_{a,a^{\prime}}}{\sum_{a^{\prime}\in\mathcal{C}_{a}^{t}}n^{t}_{a,a^{\prime}}}\bar{x}_{a,a^{\prime}}^{t}$
We observe that both in terms of memory and per-time computations, the complexity of ColME is linear, i.e., $\Theta(\mathcal{|A|})$, since every agent $a$ must keep in memory all local estimates $\bar{x}_{a,a^{\prime}}$ and number of samples $n^{t}_{a,a^{\prime}}$. And, at every $t$, for every $a^{\prime}$, it has to recompute $d^{t}_{\gamma}(a,a^{\prime})$, using the last available value for $\bar{x}_{a,a}$ (and $n^{t}_{a,a}$). Indeed, note that the local value of $\bar{x}_{a,a}$ is updated at every $t$ and this triggers a change in the optimistic distance for every $t$.
3.1 ColME’s Guarantees
We provide the the two main theoretical results from Asadi et al. [2022] which were proved only when the distributions $\{D_{a},a\in\mathcal{A}\}$ are sub-Gaussians [^2] A distribution $D$ is sub-Gaussian with parameter $\sigma^{2}$ if $\forall\lambda\in\mathbb{R}$, $\log\mathbb{E}_{x\sim D}\exp(\lambda(x-\mu))\leqslant\frac{1}{2}\lambda^{2}\sigma^{2}$..
Select the amplitude of the confidence intervals as:

### v4

@@ Abstract
In numerous settings, agents lack sufficient data to learn a model directly. Collaborating with other agents may help, but introduces a bias-variance trade-off when local data distributions differ. A key challenge is for each agent to identify clients with similar distributions while learning the model, a problem that remains largely unresolved. This study focuses on a particular instance of the overarching problem, where each agent collects samples from a real-valued distribution over time to estimate its mean. Existing algorithms face impractical per-agent space and time complexities (linear in the number of agents $|\mathcal{A}|$). To address scalability challenges, we propose a framework where agents self-organize into a graph, allowing each agent to communicate with only a selected number of peers $r$. We propose two collaborative mean estimation algorithms: one employs a consensus-based approach, while the other uses a message-passing scheme, with complexity $\mathcal{O}(r)$ and $\mathcal{O}(r\cdot\log|\mathcal{A}|)$, respectively. We establish conditions for both algorithms to yield asymptotically optimal estimates and we provide a theoretical characterization of their performance.
Codehttps://github.com/Franco-Galante/scalable-decentralized-algorithms-AAAI25
@@ 1 Introduction
Users’ devices have become increasingly sophisticated and generate vast amounts of data. This wealth of data has enabled the development of accurate and complex models. However, it has also introduced challenges related to security, privacy, real-time processing, and resource management. In response, Federated Learning (FL) has emerged as a key privacy-preserving approach for collaborative model training (Kairouz et al. 2021; Li et al. 2020). While traditional FL methods aim to develop a single model for all clients, the statistical diversity of clients’ datasets has led to the development of personalized models, designed to better align with the data distributions of individual clients (e.g., Ghosh et al. 2020; Fallah, Mokhtari, and Ozdaglar 2020; Li et al. 2021; Marfoq et al. 2021; Ding and Wang 2022).
@@
Many personalized FL strategies group clients into clusters and then tailor a model for each cluster (e.g., Ghosh et al. 2020; Sattler, Müller, and Samek 2021; Ding and Wang 2022). Ideally, clustering would group clients with similar local optimal models. However, since the optimal models are unknown a priori, model learning and cluster identification become deeply interconnected tasks. Various studies have suggested empirical measures of similarity as a workaround (e.g., Ghosh et al. 2020; Sattler, Müller, and Samek 2021), while others rely on presumed knowledge of distances across data distributions (e.g. Ding and Wang 2022; Even, Massoulié, and Scaman 2022). Nonetheless, accurately estimating these distances, especially within an FL framework where clients may possess limited data, proves to be particularly challenging. These estimation difficulties are well documented in the literature—e.g., by Even, Massoulié, and Scaman (2022) [Sec. 6]—highlighting the problem of identifying similar clients for collaborative model learning as a significant yet unresolved issue.
@@
In this paper, we focus on a fundamental aspect of the broader challenge: estimating the mean of an $\mathbb{R}^{K}$-valued distribution. This problem is often regarded as the archetypal federated learning problem (Dorner et al. 2024; Tsoy et al. 2024; Grimberg et al. 2021), but it also holds significant practical relevance across various fields, such as smart agriculture, grid management, and healthcare, where multiple sensors collect private, noisy data on identical or related variables (Adi et al. 2020).
@@
We consider an online, decentralized scenario where, at each time slot, clients receive new samples and exchange information with a limited number of peers. To the best of our knowledge, the state-of-the-art method in this setting is the Collaborative Mean Estimation algorithm (ColME) by Asadi et al. (2022). Unfortunately, ColME faces scalability issues in large systems, as both its per-agent space and time complexities are linear in the number of clients $|\mathcal{A}|$. Moreover, in its current form, ColME is applicable only to scalar mean estimation problems ($K=1$) and its convergence guarantees only hold for sub-Gaussian data distributions.
@@
We extend the methodology proposed by Asadi et al. (2022) to accommodate multidimensional data drawn from the broader class of distributions with bounded fourth moment. To address ColME’s scalability challenges, we propose that clients self-organize into a network where each client communicates with at most $r$ neighbors. Over time, this set of neighbors is pruned as clients progressively exclude the less similar ones. In this framework, we introduce two collaborative mean estimation algorithms: one based on consensus and the other on a message-passing scheme. The complexities of these algorithms are $\mathcal{O}(r)$ and $\mathcal{O}(r\cdot\log|\mathcal{A}|)$, respectively. We demonstrate that, despite each client exchanging information with only $r\ll|\mathcal{A}|$ neighbors, it is possible to achieve a convergence speedup for mean estimates by a factor of $\Omega(|\mathcal{A}|^{1/2-\phi})$, where $\phi$ can be made arbitrarily close to 00.
Lastly, we conduct preliminary experiments demonstrating how our algorithms can be adapted to federatedly learn more general machine learning models.
@@ 2 Related Work
For an overview of personalized federated learning, see the recent survey by Tan et al. (2023). Here, we highlight only the most relevant approaches related to this paper.
@@
Ghosh et al. (2020) and Sattler, Müller, and Samek (2021) were the first to propose clustered FL algorithms, which divide the clients based on the similarity of their data distributions. Similarity is empirically evaluated by the Euclidean distance between local models and by the cosine similarity of their updates. Ding and Wang (2022) study more sophisticated clustering algorithms assuming that clients can efficiently estimate some specific (pseudo)-distances across local distributions (i.e., the integral probability metrics).
@@
Beaussart et al. (2021), Chayti et al. (2022), Grimberg et al. (2021), and Even, Massoulié, and Scaman (2022) consider decentralized approaches, which allow each client to learn a personal model relying on a specific convex combination of information (gradients) from other clients. In particular, Even, Massoulié, and Scaman (2022) prove that collaboration can at most speed up the convergence time linearly in the number of similar agents and provide algorithms, which, under a priori knowledge of pairwise client distributions’ distances, achieve such speedup. The authors recognize the complexity of estimating these distances and provide practical estimation algorithms for linear regression problems, which asymptotically achieve the same speedup, scaling the number of clients but maintaining the number of clusters fixed. The generalization properties of personalized models obtained by convex combinations of clients’ models are studied in Mansour et al. (2020), while Donahue and Kleinberg (2021) look at the problem through the lens of game theory. The work most similar to ours is Asadi et al. (2022), which we describe in detail in the next section.
3 Model and Background
Table 1 lists the most important symbols used throughout the paper. Superscripts are added to variables to indicate whether they pertain to C-ColME (C), B-ColME (B), or both approaches (D).
@@ 3 Model and Background
We consider a set $\mathcal{A}$ of agents (computational units). At each time instant $t$, an agent $a\in\mathcal{A}=\{1,2,\cdots,|\mathcal{A}|\}$ generates a new sample $\bm{x}_{a}^{t}\in\mathbb{R}^{K}$, with $K\in\mathbb{N}$, drawn i.i.d. from a distribution $D_{a}$ with expected value $\bm{\mu}_{a}=\mathbb{E}[\bm{x}_{a}^{t}]$. Expected values are not necessarily distinct across agents. Indeed, given two agents $a$ and $a^{\prime}$, and a norm |||||$\cdot$||| | ⋅ | | in $\mathbb{R}^{K}$, we denote the gap between the agents’ true means by $\Delta_{a,a^{\prime}}:=||\bm{\mu}_{a}-\bm{\mu}_{a^{\prime}}||$. Let $\mathcal{C}_{a}$ be the group of agents with the same true mean as $a$, (i.e., those for which $\Delta_{a,a^{\prime}}=0$). In the following, we will refer to $\mathcal{C}_{a}$ as the ‘similarity class of $a$.’
@@
The goal of each agent $a\in\mathcal{A}$ is to estimate its mean $\bm{\mu}_{a}$. To this end, at each time $t$ the agent can compute its local mean estimate $\bar{\bm{x}}^{t}_{a,a}=\frac{1}{t}\sum_{t^{\prime}=1}^{t}\bm{x}^{t^{\prime}}_{a}$ over the $t$ available samples. Additionally, the agent can obtain a more accurate estimate by leveraging information from other agents in $\mathcal{A}$, provided it can identify those who share the same true mean.
For the scalar case, (i.e., when $K=1$) Asadi et al. (2022) proposed ColME as a collaborative algorithm for mean estimation. It relies on two key steps, executed concurrently: i) the identification of the similarity classes; ii) the collaboration with agents believed to belong to the same class to improve the local estimate. To ensure direct comparability with the results in (Asadi et al. 2022) and simplify notation, in what follows, we focus on the scalar case. Readers interested in the general case can find the analysis in Appendix C. All the appendices referred to in this work are accessible at (Galante, Neglia, and Leonardi 2024).
Symbol Description
$\mathcal{A}$ Set of agents (computational units)
$D_{a}$ Distribution of agent $a\in\mathcal{A}$
$\bm{\mu}_{a}$ True mean of distribution $D_{a}$, i.e., $\bm{\mu}_{a}=\mathbb{E}[\bm{x}_{a}^{t}]$
$\bar{\bm{x}}^{t}_{a}$ Vectorial sample $\bar{\bm{x}}^{t}_{a}\in\mathbb{R}^{K}$ drawn from $D_{a}$
$\Delta_{a,a^{\prime}}$ Gap between agents $a$ and $a^{\prime}$ true means
$\mathcal{C}_{a}$ Similarity class of $a$, $\mathcal{C}_{a}=\{a^{\prime}\in\mathcal{A}|\Delta_{a,a^{\prime}}=0\}$
$\mathcal{C}_{a}^{t}$ Estimated similarity class of $a$ at time $t$
$\bar{x}^{t}_{a,a}$ Local mean estimate of agent $a$ at $t$
$\bar{x}^{t}_{a,a^{\prime}}$ Local mean estimate of agent $a^{\prime}$ by agent $a$ at $t$
$n^{t}_{a,a^{\prime}}$ Number of samples used to compute $\bar{x}^{t}_{a,a^{\prime}}$ at $t$
$\hat{\mu}_{a}^{t}$ Collaborative mean estimate at $t$
$\beta_{\gamma}(n)$ Width of the confidence interval
$d_{\gamma}^{t}(a,a^{\prime})$ Optimistic distance between agents $a$ and $a^{\prime}$
$\zeta_{a}$ Time to identify same-class neighbors w.h.p.
$\tau_{a}$ Time to obtain $(\epsilon,\delta)$ convergence
$\mathcal{G}$ Collaborative graph $\mathcal{G}(\mathcal{A},\mathcal{E})$
$\mathcal{G}^{t}$ Pruned collaborative graph $\mathcal{G}^{t}(\mathcal{A},\mathcal{E}^{t})$
$\mathcal{N}_{a}$ Neighborhood of agent $a$ (or up to distance $d$: $\mathcal{N}_{a}^{d}$)
$r$ Upper bound on $a$’s neighborhood size $|\mathcal{N}_{a}|$
$\mathcal{CC}_{a}$ Agents in the connected component of the subgraph
induced by agents in $\mathcal{C}_{a}$ to which agent $a$ belongs
$\mathcal{CC}_{a}^{d}$ Same as $\mathcal{CC}_{a}$ but with agents up to $d$-hops from $a$
Table 1: Notation Summary
@@
[…]
$d^{t}_{\gamma}(a,a^{\prime}):=|\bar{x}_{a,a}-\bar{x}_{a,a^{\prime}}|-\beta_{\gamma}(t)-\beta_{\gamma}(n^{t}_{a,a^{\prime}})\leqslant 0.$ (1)
To achieve the most accurate estimation of its similarity class, agent $a$ could retrieve from each other peer $a^{\prime}$ the most recent estimate $\bar{x}^{t}_{a,a^{\prime}}$ at each time $t$. However, this approach would result in a per-agent communication burden of $\mathcal{O}(|\mathcal{A}|)$. To mitigate this effect, node $a$ cyclically queries a single node from $\mathcal{C}_{a}^{t-1}$ at each time instant $t$, according to a Round-Robin scheme. This leads to the following update rule for $\mathcal{C}_{a}^{t}$:
@@
and we observe that by construction $\mathcal{C}_{a}^{t}\subseteq\mathcal{C}_{a}^{t-1}$. Initially, the agent sets $\mathcal{C}_{a}^{0}=\mathcal{A}$ and progressively makes irreversible decisions to remove agents that are deemed too dissimilar.
@@
Each node $a$ computes an estimate $\hat{\mu}_{a}^{t}$ of its true mean $\mu_{a}$ combining the available estimates according to a simple weighting scheme, where the number of samples $n^{t}_{a,a^{\prime}}$ are the weights:
$\hat{\mu}_{a}^{t}=\sum_{a^{\prime}\in\mathcal{C}_{a}^{t}}\frac{n^{t}_{a,a^{\prime}}}{\sum_{a^{\prime}\in\mathcal{C}_{a}^{t}}n^{t}_{a,a^{\prime}}}\bar{x}_{a,a^{\prime}}^{t}.$
ColME’s theoretical guarantees.
Asadi et al. (2022) prove that if data distributions $\{D_{a}\}_{a\in\mathcal{A}}$ are sub-Gaussian and the amplitude of the confidence intervals is selected as:
$\beta_{\gamma}(n)=\sigma\sqrt{\frac{2}{n}\left(1+\frac{1}{n}\right)\ln(\sqrt{(n+1)}/\gamma)},$ (3)
then, with high probability, client $a$ correctly identifies its similarity class ($\mathcal{C}_{a}^{t}=\mathcal{C}_{a}$) and obtains an $\epsilon$-accurate estimate for $t$ larger than two opportune constants $\zeta_{a}$ and $\tau_{a}$, respectively (see Appendix A for the expressions and the results’ statements).
ColME’s limitations.
Asadi et al. (2022) acknowledge in their paper the main limitations of ColME: each agent requires a memory footprint and computational complexity proportional to $|\mathcal{A}|$. Indeed, each agent $a$ must store all neighbors’ local estimates $\bar{x}^{t}_{a,a^{\prime}}$ and the corresponding sample counts $n^{t}_{a,a^{\prime}}$. Additionally, as agent $a$ receives a new sample at each time $t$, it updates its local estimate $\bar{x}^{t}_{a,a}$ (and also $n^{t}_{a,a}$). This update affects the distance $d_{\gamma}^{t}(a,a^{\prime})$, which thus must be recomputed for all $a^{\prime}\in\mathcal{C}_{a}^{t}$. This leads to a per-agent time and space complexity of $\mathcal{O}(|\mathcal{A}|)$ per time slot, which becomes impractical in large-scale systems. Moreover, while the Round-Robin query scheme reduces the communication burden, it introduces a significant delay in the estimation as $\zeta_{a}\in\mathcal{O}(|\mathcal{A}|)$.
4 Scalable Algorithms over a Graph $\mathcal{G}$
Online mean estimation can be made $\tilde{\mathcal{O}}(1)$ with our scalable approaches: C-ColME (Sec. 4.1), and B-ColME (Sec. 4.2). Both algorithms consider agents $\mathcal{A}$ organized in a fixed graph $\mathcal{G}(\mathcal{A},\mathcal{E})$ and restrict communication to pairs of agents adjacent in $\mathcal{G}$. Let $\mathcal{N}_{a}$ and $\mathcal{N}_{a}^{d}$ denote the set of neighbors of agent $a$ and the set of agents at distance at most $d$ from $a$, respectively. Let $r$ represent the maximum size of any agent’s neighborhood in $\mathcal{G}$, i.e., $|\mathcal{N}_{a}|\leq r$ for all $a\in\mathcal{A}$.
Consider the subgraph $\mathcal{G}^{\prime}$ of $\mathcal{G}$ induced by the agents in $\mathcal{C}_{a}$. Let $\mathcal{CC}_{a}$ denote the (initially unknown) set of agents in the connected component of $\mathcal{G}^{\prime}$ to which $a$ belongs, and let $\mathcal{CC}_{a}^{d}\subset\mathcal{CC}_{a}$ represent the subset of agents within $\mathcal{CC}_{a}$ that are at most $d$ hops away from $a$.
Each agent $a$ aims to identify which nodes in its neighborhood $\mathcal{N}_{a}$ belong to its similarity class $\mathcal{C}_{a}$. To achieve this, agent $a$ receives at time $t$ an updated local mean estimate $\bar{x}^{t^{\prime}}_{a,a^{\prime}}$ from each neighbor $a^{\prime}\in\mathcal{N}_{a}$. We denote with $\mathcal{C}_{a}^{t}\subseteq\mathcal{N}_{a}$ the set of neighbors that agent $a$ deems to belong to its own similarity class at time $t$, initially: $\mathcal{C}_{a}^{0}=\mathcal{N}_{a}$. Similarly to ColME, at each time $t$, agent $a$ first computes the distance $d_{\gamma}^{t}(a,a^{\prime})$ for every $a^{\prime}\in\mathcal{C}_{a}^{t-1}$ according to (1) and then updates $\mathcal{C}_{a}^{t}$ according to (2). As for ColME, $\mathcal{C}_{a}^{t}\subseteq\mathcal{C}_{a}^{t-1}$ and as soon as $a$ removes $a^{\prime}$ from $\mathcal{C}_{a}^{t}$, it stops communicating with $a^{\prime}$. Subsequently, communication occur over the pruned graph $\mathcal{G}^{t}=(\mathcal{A},\mathcal{E}^{t})$, where $\mathcal{E}^{t}=\{(a,a^{\prime})\in\mathcal{E}:a^{\prime}\in~{}\mathcal{C}_{a}^{t}\}$.
The theoretical guarantees of our algorithms hold under more general settings than those in (Asadi et al. 2022). In particular, they apply to any set of distributions $\{D_{a}\}_{a\in\mathcal{A}}$ for which the following assumption is satisfied:
Assumption 1.
There exists a positive function $\beta_{\gamma}(\cdot)\in o(1)$ such that the true mean belongs to all intervals centered in $\bar{x}_{a,a}^{t}$ of width ±$\pm\beta_{\gamma}(t)$ for $t\in\mathbb{N}$ with confidence $1-2\gamma$, namely:
Assumption 1 is satisfied by sub-Gaussian distributions (SGD) with parameter $\sigma^{2}$, by selecting $\beta_{\gamma}(\cdot)$ as in (3). In Appendix B, we show that the assumption also holds for bounded fourth-moment distributions (BFMD) for $\beta_{\gamma}(\cdot)$ chosen as follows:
$\beta_{\gamma}(n)=\left(2\frac{(\kappa+3)\sigma^{4}}{\gamma}\right)^{\frac{1}{4}}\left(\frac{1+\ln^{2}n}{n}\right)^{\frac{1}{4}},$ (5)
where $\sigma^{2}$ bounds the variance of the distributions $\{D_{a},\forall a\in\mathcal{A}\}$ and $\kappa\sigma^{4}$ their fourth moment. When all the variables are identically distributed, $\kappa$ corresponds to the kurtosis. Moreover, for distributions with a larger number of bounded moments, tighter expressions can be derived for $\beta_{\gamma}(\cdot)$ (see Remark 1 in Appendix B). In what follows, we assume that Assumption 1 is always satisfied.
We aim first to determine the time needed for all agents in the connected component $\mathcal{CC}_{a}$ to identify the subset of neighbors residing in their similarity class, i.e., $\mathcal{C}_{a^{\prime}}^{t}=\mathcal{C}_{a^{\prime}}\cap\mathcal{N}_{a^{\prime}}$, $\forall a^{\prime}\in\mathcal{CC}_{a}$. Following a similar approach to Asadi et al. (2022, Theorem 1) we can prove that: